(a+b)^2

(a+b)^2

Grade:6

5 Answers

charan
8 years ago
(a^2+b^2+2ab
 
HARINATH
8 years ago
(a+b)2 = a2+2ab+b2
SHANMUKESHWAR
8 years ago
+2ab+b22= a2 (a+b)
RAJU
8 years ago
(a-b)2+4ab
OUTPASS
8 years ago
(a+b)= a2+2ab+b2